Problem 5-57 Spreadsheet Problem (LG5-2, LG5-9) Consider a person who begins contributing to a retirement plan at age 25 and pls be simple and clear with the answer and explanation.
Answer #1
Solution 1
Here we have 4 annuities of 10 years each. We will calculate the amount of each annuity and will keep the same invested till age is 65
FV of annuity
The formula for the future value of an ordinary annuity, as opposed to an annuity due, is as follows:
P = PMT x ((((1 + r) ^ n) - 1) / r)
Where:
P = the future value of an annuity stream To be computed
PMT = the dollar amount of each annuity payment Given
r = the effective interest rate (also known as the discount rate) 10%
n = the number of periods in which payments will be made 10 years
Annuity value factor =((((1 + 10%) ^ 10) - 1) / 10%)
Annuity value factor                                              15.9374
Annuity 1 2 3 4
Annuity amount $                                              3,700 $                             5,700 $                              10,700 $                           15,700
Interest rate 10% 10% 10% 10%
Amount remained invested for next years 30 20 10 0
Maturity value of each annuity= Annuity amount * Annuity factor
Maturity value of each annuity= 3700*15.9374 5700*15.9374 10700*15.9374 15700*15.9374
Maturity value of each annuity= $                                      58,968.38 $                     90,843.18 $                      170,530.18 $                   250,217.18
Maturity value of each annuity at age 65= =58968.38*(1+10%)^30 =90843.18*(1+10%)^20 =170530.18*(1+10%)^10 =250217.18*(1+10%)^0
Maturity value of each annuity at age 65= $                                 1,028,962.98 $                   611,147.49 $                      442,311.37 $                   250,217.18
Total fund value =1028962.98+611147.49+442311.37+250217.18
Total fund value $                                 2,332,639.02
Solution 2
Now this corpus is converted into an annuity so the PV of all annuity payments should be equal to 2,332,639.02
PV of annuity
P = PMT x (((1-(1 + r) ^- n)) / r)
Where:
P = the present value of an annuity stream $                                 2,332,639.02
PMT = the dollar amount of each annuity payment To be calculated
r = the effective interest rate (also known as the discount rate) 9%
n = the number of periods in which payments will be made 40
PV of annuity during retirement= PMT x (((1-(1 + r) ^- n)) / r)
2332639.021= PMT * (((1-(1 + 9%) ^- 40)) / 9%)
2332639.021= PMT * 10.757360195239
Each annuity payment= 2332639.021/10.757360195239
Each annuity payment= $                                    216,841.21